# AI Research Agents for Machine Learning: Search, Exploration, and Generalization in MLE-bench

### v1

@@ 1 Introduction
Science is based on searching the open-ended space of hypotheses and testing them in a controlled experiment (Langley et al., 1987). Recent breakthroughs have resulted in artificial intelligence (AI) agents that offer great potential to automate the scientific discovery process (Yamada et al., 2025; Swanson et al., 2024; Boiko et al., 2023). A key obstacle to improving research agents is that their designs entangle several factors for performance, making it difficult to pinpoint sources of improvement via controlled experiments at scale. These factors span algorithm design, concrete implementation, and ability to leverage compute, as performance gains accrue only if no layer in the stack bottlenecks the benefits from additional compute resources (Fig. 7). This challenge is exemplified in MLE-bench (Chan et al., 2025), a benchmark where AI agents compete in Kaggle competitions to solve real-world machine learning (ML) problems. Notably, the state-of-the-art approach AIDE (Jiang et al., 2025) does not fully disentangle these performance factors, providing limited insight into which components primarily drive the agent’s performance and where improvements are needed.
Figure 2: Overview of AIRA. Given a problem specification, AIRA maintains a search graph whose nodes are (partial) solutions. At each iteration, the agent (1) selects nodes via a selection policy, (2) picks an operator via an operator policy and applies this operator to the node, and (3) scores the resulting solution via a fitness function. Here, a greedy node-selection strategy applies the improve operator to the highest-scoring node.
We start by formalizing the design of AI research agents as a search algorithm composed of two components: The first one is the search policy which is used to navigate the space of candidate solutions, and the second is the operators which iteratively modifies existing solutions to generate new candidate solutions. In this framework (Fig. 2), AIDE is represented as a greedy search algorithm that, at each step, applies one of its code operators (i.e., Draft, Debug, Improve) to the current best solution. This allows us to disentangle the effect of the search from that of the operators.
To assess alternative search algorithms, we develop sophisticated agents performing Evolutionary and Monte Carlo Tree Search (MCTS) (De Jong, 2017; Kocsis and Szepesvári, 2006). We empirically show that AIDE’s operators, rather than the search algorithm, are a bottleneck to better performance.
@@ 2.1 Graph–based Search Framework
Definition 1 (Components of the Search Algorithm). A graph-based search algorithm is specified by the tuple $\bigl{(}\mathcal{F},\pi_{\text{sel}},\mathcal{O},\pi_{\text{op}},\tau\bigr{)}$:
@@
At each iteration, the agent selects existing promising artifacts, applies transformation operators, and updates the search graph, propelling the discovery process forward. We discuss specific instantiations of search algorithms in Sections 2.34.2. In all our instantiations: 1 the termination criterion is set as the wall-clock time or the maximum number of artifacts, whichever happens first; 2 the fitness function $\mathcal{F}$ is defined for each node by the operator that generates or modifies it (i.e., $\mathcal{F}$ is not global); 3 all operators are LLM-driven except for the Memory operator, which is defined by hand.
@@ 2.2 Operators
Section 3 contains further implementation details.
@@ 2.3 Re-casting AIDE in our Notation
AIDE (Jiang et al., 2025) is an LLM-driven agent that frames problem-solving as a tree-search over Python scripts. In the tuple $(\mathcal{F},\pi_{\text{sel}},\mathcal{O},\pi_{\text{op}},\tau)$ from Section 2.1 its components are:
@@
This combination of $\pi_{\text{sel}},\pi_{\text{op}}$, and $\mathcal{O}_{\text{AIDE}}$ defines the baseline agent we denote AIDEgreedy (also summarized in Section 9). When the agent uses an alternative search policy while keeping $\mathcal{O}_{\text{AIDE}}$ unchanged, we write AIDEsearch_policy to identify change. For example, AIDEmcts uses the AIDE operator set and MCTS as the search policy.
@@ 3.1 AIRA-dojo
An agent’s environment greatly influences its performance. To systematically study the space of agentic policies (see Fig. 7), we introduce the AI Research Agent (AIRA) dojo—a scalable and customizable framework for AI research agents. AIRA-dojo provides the environment in which agents operate, along with abstractions for operators and policies as described in Section 2, and tasks that define evaluation criteria for agent performance. Using these abstractions, we implement and evaluate four search policies: MCTS, Evolutionary, Greedy, and our own implementation of AIDE. We hope that AIRA-dojo’s scalability and customizability, together with the provided agent and task implementations, will support and advance future research in the community.
The infrastructure design of AIRA-dojo was informed by several reliability and performance constraints, as discussed in more detail in Section 13.
@@ 3.3 MLE-bench
Our experiments showed that existing methods exhibit high variance in this benchmark (see Section 15). Focusing on a smaller subset of tasks allows us to allocate more seeds per task and increase our confidence in our results.
@@ 3.4 Experimental Details
Environment. Each candidate agent is launched in a freshly initialized, sandboxed process in AIRA-dojo. This guarantees that file systems and environment variables are isolated across evaluations, and there is no cross-agent interference. Every sandbox is provisioned with a fixed hardware quota: 1 dedicated H200 GPU, 24 logical CPU cores, 100 GB of RAM, and 1 TB of additional scratch storage. Internet access is permitted solely for fetching third-party packages and model checkpoints.
@@
See Section 11 for further implementation details.
@@ 4.1 Operators $\mathcal{O}_{\text{AIRA}}$
    Prompt-adaptive complexity. We introduce a dynamic complexity cue within the system prompt in order to guide the complexity of artifacts generated by the draft and improve operators. The complexity is determined by the number of children, $n_{\mathrm{c}}$, of the node being processed:
    $\mathrm{complexity}(n_{\mathrm{c}})=\text{``minimal''}\leavevmode\nobreak\ \mathrm{if}\leavevmode\nobreak\ n_{\mathrm{c}}<2,\leavevmode\nobreak\ \text{``moderate''}\leavevmode\nobreak\ \mathrm{if}\leavevmode\nobreak\ 2\leq n_{\mathrm{c}}<4,\leavevmode\nobreak\ \text{``advanced''}\leavevmode\nobreak\ \mathrm{if}\leavevmode\nobreak\ n_{\mathrm{c}}\geq 5$
@@
    Think Tokens. For reasoning models, we use the operators’ system prompts to explicitly encourage them to use thinking tokens for structured reasoning and reflection. These thoughts are stripped from the final answer—remaining invisible to other operators (such as memory). On average, we observe a 2×2$\times$2 × increase in completion tokens generated by the AIRA operator set (see Section 12).
@@ 4.2 Agents
In this section, we introduce three agents that combine the operators $\mathcal{O}_{\text{AIRA}}$—proposed in Section 4.1—with distinct search policies (see Appendix 14 for the rationale behind their selection). Each agent uses the same proxy–fitness function $\mathcal{F}$ (5-fold CV) and the same termination criterion $\tau$ (based on wall-clock time or artifact cap).
AIRAgreedy. This agent employs greedy search (Section 2.3) using the $\mathcal{O}_{\text{AIRA}}$ operator set. Any performance improvement of AIRAgreedy over AIDEgreedy directly reflects the benefit of the new operators, since the only difference between the two is the operator set.
AIRAmcts. This agent uses Monte-Carlo Tree Search (MCTS) Kocsis and Szepesvári (2006) with $\mathcal{O}_{\text{AIRA}}$ operator set. Our implementation of MCTS follows the canonical loop (selection, expansion, evaluation, and backup), but omits simulated roll-outs: the leaf value of expanded nodes is given directly by the proxy fitness function $\mathcal{F}$:
@@ 5.1 Analyzing the Performance of the Current SoTA
What is the effect of memory? To quantify the impact of the memory operator, we conduct a controlled ablation comparing the performance of AIDE with and without the memory operator enabled. As shown in Fig. 3, both variants achieve nearly identical mean medal rates. This suggests that memory is not a driving factor behind AIDE’s strong performance.
@@
What is the effect of exploration at the search level with AIDE’s operators? AIDE uses a selection policy that does not explore and always greedily selects the most promising candidate. To evaluate the impact of the search policy as a modulator of the exploration-exploitation tradeoff when searching using AIDE’s operators, we replaced the operators in AIRAmcts with that of AIDE and varied the UCT exploration constant $c_{\mathrm{UCT}}\in\{0,0.25,0.75\}$. The $c_{\mathrm{UCT}}$, which modulates the degree of exploration in MCTS, allows us to isolate the effect of search-level exploration on downstream performance. The resulting medal rates (Fig. 3) showed only marginal differences across all $c_{\mathrm{UCT}}$ values, supporting our hypothesis that search-level exploration is constrained by the interaction between the operator set and the search policy. We observe the same limitation when performing the same operator replacement in AIRAevo.
What is the performance profile of AIDE? To examine the improvement gains over time, we plot AIDE’s anytime performance, which is the average medal rate achieved by the agent if the search was to terminate at time $t$. The results, summarized in Fig. 4, suggest that while the agent’s perceived performance (validation score) continues to improve over time, the true test performance plateaus, even slightly decreases over time. These findings suggest that overfitting might be a fundamental limitation to the agent’s performance. We further investigate this in Section 5.3.
@@ 5.2 AIRA Beyond Greedy
Based on the observation that search policies yield no performance gain with AIDE’s operators, this section is divided into two parts: 1 assessing the effectiveness of our improved operators (see Section 4.1), 2 examining their interplay with the more advanced search policies. The results are summarized in Fig. 5, with a detailed breakdown of performance for each task in Section 17.
@@
Evaluating search policies. Equipped with an improved set of operators, we now explore whether combining them with advanced search methods can further enhance performance. The results (Fig. 5) show that for R1, AIRAmcts achieves the best performance, achieving state-of-the-art performance on MLE-bench Lite with an average medal rate of 47%. Both AIRAmcts and AIRAgreedy achieve silver-or-above medal rates ($\approx$ 36.5%), outperforming the baseline by 4.5 absolute points. In terms of gold medals, AIRAgreedy performs best, reaching 27%. For o3, AIRAmcts and AIRAgreedy perform comparably in terms of any medal and silver-or-above medal rates, while AIRAmcts outperforms AIRAgreedy in gold medals by 2.2 percentage points. Although agents using R1 and o3 achieve similar overall medal rates, those using o3 perform better when considering gold and silver-or-above medal rates. Specifically, AIRAmcts increases the gold medal rate from 25.68 % with R1 to 30.91 % with o3, a relative increase of 20 %.
Collectively, what stands out is that all search policies combined with AIRA operators outperform AIDEgreedy. Furthermore, the rankings between agents using the AIRA operators with different search strategies differs significantly from that observed in Section 5.1.
Finally, Fig. 4 shows the anytime performance of each agent over the 24-hour window defined by MLE-bench. For conciseness, we will focus our discussion on performance with R1, as the results with o3 follow a similar pattern. We observe that the rankings between agents change over time. For example, at the 3-hour mark, the performance gap between AIRAgreedy and AIRAmcts is notable. This gap narrows by the 10-hour mark, where all agents perform similarly. Meaningful divergence in performance appears only after 19 hours. These trends reflect the interaction between rankings and the resources provided, as shown in Figure 7.
@@ 5.3 The Generalization Gap: Searching with a Proxy Evaluation
Figure 6: a) The validation–test metric mismatch. Results shown are for agents using R1. Bars depict the absolute performance gap between three configurations: (i) using the validation metric for both intermediate (search) and final (submission) node selection; (ii) using the validation metric only for search and the test metric for selection; and (iii) using the test metric for both search and selection. b) Bridging the validation–test gap. Medal rate achieved by two final node selection strategies as a function of $k$: (i) randomly sampling $k$ nodes and reporting the highest test score among them; (ii) selecting the top $k$ nodes by validation score and reporting the highest test score among those. As $k$ increases, the validation-based strategy closes the gap to the upper-bound performance given by the best test score over the entire search graph.
@@
How large is the impact of the generalization gap on performance? We begin by comparing two extremes: Val/Val—both intermediate and final node selection use the validation score (standard practice); Test/Test—an oracle baseline using the true test score at both stages. As shown in Fig. 6, searching based on the test score instead of the validation score improves performance by 9.4 % and 12.4 % for AIRAmcts and AIRAevo. The gap is even higher for the agents using a greedy search policy, reaching 15 % for AIRAgreedy and 16.6 % for AIDEgreedy. For the rest of this section, we investigate the nature of this gap in performance and how to close it.
How much of the gap can be attributed to final node selection? To answer this question, we consider the following two settings: Val/Val; and Val/Test—the intermediate node selection uses the validation score, while the final node selection is done based on the test scores. The results (Fig. 6) indicate that by selecting the best node in a search graph constructed without privileged access, i.e., based on validation signal, all agents achieve a performance boost of 9 to 11.6 absolute points. Crucially, comparing Val/Test with Test/Test shows that oracle final-node selection eliminates the gap between the standard (Val/Val) and oracle (Test/Test) settings for both AIRAmcts and AIRAevo. Even for the agents implementing the simpler greedy policy, selecting the best node in the final search graph closes more than 60 % of the gap. These findings highlight robust final-node-selection strategies as a promising path to higher performance.
Bridging the gap through multiple submissions. A straightforward remedy is to hedge against validation score noise by selecting multiple promising nodes. Specifically, among the top-$k$ nodes ranked by validation score in the search graph, we report the test score of the best-performing one. As a baseline, we repeat the process for a set of random-$k$ nodes, and summarize the results in Fig. 6. We highlight the following observations: 1 the larger gap between top-k and random-k in agents using non-greedy algorithms, AIRAevo and AIRAmcts, suggest greater diversity in their search graphs; 2 with as little as 3 top-k submissions agents achieve an additional 10 % of performance; 3 the top validation nodes are informative of the best performing nodes.

### v2

@@ 1 Introduction
Science is based on searching the open-ended space of hypotheses and testing them in a controlled experiment (Langley et al., 1987). Recent breakthroughs have resulted in artificial intelligence (AI) agents that offer great potential to automate the scientific discovery process (Yamada et al., 2025; Swanson et al., 2024; Boiko et al., 2023). A key obstacle to improving research agents is that their designs entangle several factors for performance, making it difficult to pinpoint sources of improvement via controlled experiments at scale. These factors span algorithm design, concrete implementation, and ability to leverage compute, as performance gains accrue only if no layer in the stack bottlenecks the benefits from additional compute resources (Fig.˜7). This challenge is exemplified in MLE-bench (Chan et al., 2025), a benchmark where AI agents compete in Kaggle competitions to solve real-world machine learning (ML) problems. Notably, the state-of-the-art approach AIDE (Jiang et al., 2025) does not fully disentangle these performance factors, providing limited insight into which components primarily drive the agent’s performance and where improvements are needed.
Figure 2: Overview of AIRA. Given a problem specification, AIRA maintains a search graph whose nodes are (partial) solutions. At each iteration, the agent (1) selects nodes via a selection policy, (2) picks an operator via an operator policy and applies this operator to the node, and (3) scores the resulting solution via a fitness function. Here, a greedy node‐selection strategy applies the improve operator to the highest‐scoring node.
We start by formalizing the design of AI research agents as a search algorithm composed of two components: The first one is the search policy which is used to navigate the space of candidate solutions, and the second is the operators which iteratively modifies existing solutions to generate new candidate solutions. In this framework (Fig.˜2), AIDE is represented as a greedy search algorithm that, at each step, applies one of its code operators (i.e., Draft, Debug, Improve) to the current best solution. This allows us to disentangle the effect of the search from that of the operators.
To assess alternative search algorithms, we develop more sophisticated agents performing Evolutionary and Monte Carlo Tree Search (MCTS) (De Jong, 2017; Kocsis and Szepesvári, 2006). We empirically show that AIDE’s operators, rather than the search algorithm, are a bottleneck to better performance.
@@
Rerunning the best-performing agents with our latest version of AIRA-dojo further increased performance to 55 % on MLE-Bench Lite. For details on these experiments and results on the complete MLE-bench suite of tasks, see Section˜11.
@@ 2.1 Graph–based Search Framework
Definition 1 (Components of the Search Algorithm). A graph‑based search algorithm is specified by the tuple $\bigl(\mathcal{F},\pi_{\text{sel}},\mathcal{O},\pi_{\text{op}},\tau\bigr)$:
@@
At each iteration, the agent selects existing promising artifacts, applies transformation operators, and updates the search graph, propelling the discovery process forward. We discuss specific instantiations of search algorithms in Sections 2.3, 4.2. In all our instantiations: 1 the termination criterion is set as the wall-clock time or the maximum number of artifacts, whichever happens first; 2 the fitness function $\mathcal{F}$ is defined for each node by the operator that generates or modifies it (i.e., $\mathcal{F}$ is not global); 3 all operators are LLM-driven except for the Memory operator, which is defined by hand.
@@ 2.2 Operators
Section˜3 contains further implementation details.
@@ 2.3 Re-casting AIDE in our Notation
AIDE (Jiang et al., 2025) is an LLM-driven agent that frames problem-solving as a tree-search over Python scripts. In the tuple $(\mathcal{F},\pi_{\text{sel}},\mathcal{O},\pi_{\text{op}},\tau)$ from Section˜2.1 its components are:
@@
[…]
@@ 3.1 AIRA-dojo
An agent’s environment greatly influences its performance. To systematically study the space of agentic policies (see Fig.˜7), we introduce the AI Research Agent (AIRA) dojo—a scalable and customizable framework for AI research agents. AIRA-dojo provides the environment in which agents operate, along with abstractions for operators and policies as described in Section˜2, and tasks that define evaluation criteria for agent performance. Using these abstractions, we implement and evaluate four search policies: MCTS, Evolutionary, Greedy, and our own implementation of AIDE. We hope that AIRA-dojo’s scalability and customizability, together with the provided agent and task implementations, will support and advance future research in the community.
The infrastructure design of AIRA-dojo was informed by several reliability and performance constraints, as discussed in more detail in Section˜14.
@@ 3.3 MLE-bench
Our experiments showed that existing methods exhibit high variance in this benchmark (see Section˜16). Focusing on a smaller subset of tasks allows us to allocate more seeds per task and increase our confidence in our results.
@@ 3.4 Experimental Details
Environment. Each candidate agent is launched in a freshly initialized, sandboxed process in AIRA-dojo. This guarantees that file systems and environment variables are isolated across evaluations, and there is no cross-agent interference. Every sandbox is provisioned with a fixed hardware quota: 1 dedicated H200 GPU, 24 logical CPU cores, 100 GB of RAM, and 1 TB of additional scratch storage. Internet access is permitted solely for fetching third‑party packages and model checkpoints.
@@
See Section˜12 for further implementation details.
@@ 4.1 Operators $\mathcal{O}_{\text{AIRA}}$
    Prompt‑adaptive complexity. We introduce a dynamic complexity cue within the system prompt in order to guide the complexity of artifacts generated by the draft and improve operators. The complexity is determined by the number of children, $n_{\mathrm{c}}$, of the node being processed:
    $\mathrm{complexity}(n_{\mathrm{c}})=\text{``minimal''}\penalty 10000\ \mathrm{if}\penalty 10000\ n_{\mathrm{c}}<2,\penalty 10000\ \text{``moderate''}\penalty 10000\ \mathrm{if}\penalty 10000\ 2\leq n_{\mathrm{c}}<4,\penalty 10000\ \text{``advanced''}\penalty 10000\ \mathrm{if}\penalty 10000\ n_{\mathrm{c}}\geq 5$
@@
    Think Tokens. For reasoning models, we use the operators’ system prompts to explicitly encourage them to use thinking tokens for structured reasoning and reflection. These thoughts are stripped from the final answer—remaining invisible to other operators (such as memory). On average, we observe a $2\times$ increase in completion tokens generated by the AIRA operator set (see Section˜13).
@@ 4.2 Agents
In this section, we introduce three agents that combine the operators $\mathcal{O}_{\text{AIRA}}$—proposed in Section˜4.1—with distinct search policies (see Appendix 15 for the rationale behind their selection). Each agent uses the same proxy–fitness function $\mathcal{F}$ (5-fold CV) and the same termination criterion $\tau$ (based on wall-clock time or artifact cap).
AIRAgreedy. This agent employs greedy search (Section˜2.3) using the $\mathcal{O}_{\text{AIRA}}$ operator set. Any performance improvement of AIRAgreedy over AIDEgreedy directly reflects the benefit of the new operators, since the only difference between the two is the operator set.
AIRAmcts. This agent uses Monte-Carlo Tree Search (MCTS) Kocsis and Szepesvári (2006) with $\mathcal{O}_{\text{AIRA}}$ operator set. Our implementation of MCTS follows the canonical loop (selection, expansion, evaluation, and backup), but omits simulated roll‑outs: the leaf value of expanded nodes is given directly by the proxy fitness function $\mathcal{F}$:
@@ 5.1 Analyzing the Performance of the Current SoTA
What is the effect of memory? To quantify the impact of the memory operator, we conduct a controlled ablation comparing the performance of AIDE with and without the memory operator enabled. As shown in Fig.˜3, both variants achieve nearly identical mean medal rates. This suggests that memory is not a driving factor behind AIDE’s strong performance.
@@
What is the effect of exploration at the search level with AIDE’s operators? AIDE uses a selection policy that does not explore and always greedily selects the most promising candidate. To evaluate the impact of the search policy as a modulator of the exploration-exploitation tradeoff when searching using AIDE’s operators, we replaced the operators in AIRAmcts with that of AIDE and varied the UCT exploration constant $c_{\mathrm{UCT}}\in\{0,0.25,0.75\}$. The $c_{\mathrm{UCT}}$, which modulates the degree of exploration in MCTS, allows us to isolate the effect of search-level exploration on downstream performance. The resulting medal rates (Fig.˜3) showed only marginal differences across all $c_{\mathrm{UCT}}$ values, supporting our hypothesis that search-level exploration is constrained by the interaction between the operator set and the search policy. We observe the same limitation when performing the same operator replacement in AIRAevo.
What is the performance profile of AIDE? To examine the improvement gains over time, we plot AIDE’s anytime performance, which is the average medal rate achieved by the agent if the search was to terminate at time $t$. The results, summarized in Fig.˜4, suggest that while the agent’s perceived performance (validation score) continues to improve over time, the true test performance plateaus, even slightly decreases over time. These findings suggest that overfitting might be a fundamental limitation to the agent’s performance. We further investigate this in Section˜5.3.
@@ 5.2 AIRA Beyond Greedy
Based on the observation that search policies yield no performance gain with AIDE’s operators, this section is divided into two parts: 1 assessing the effectiveness of our improved operators (see Section˜4.1), 2 examining their interplay with the more advanced search policies. The results are summarized in Fig.˜5, with a detailed breakdown of performance for each task in Section˜18.
@@
Evaluating search policies. Equipped with an improved set of operators, we now explore whether combining them with advanced search methods can further enhance performance. The results (Fig.˜5) show that for R1, AIRAmcts achieves the best performance, achieving state-of-the-art performance on MLE-bench Lite with an average medal rate of 47%. Both AIRAmcts and AIRAgreedy achieve silver-or-above medal rates ($\approx$ 36.5%), outperforming the baseline by 4.5 absolute points. In terms of gold medals, AIRAgreedy performs best, reaching 27%. For o3, AIRAmcts and AIRAgreedy perform comparably in terms of any medal and silver-or-above medal rates, while AIRAmcts outperforms AIRAgreedy in gold medals by 2.2 percentage points. Although agents using R1 and o3 achieve similar overall medal rates, those using o3 perform better when considering gold and silver-or-above medal rates. Specifically, AIRAmcts increases the gold medal rate from 25.68 % with R1 to 30.91 % with o3, a relative increase of 20 %.
Collectively, what stands out is that all search policies combined with AIRA operators outperform AIDEgreedy. Furthermore, the rankings between agents using the AIRA operators with different search strategies differs significantly from that observed in Section˜5.1.
Finally, Fig.˜4 shows the anytime performance of each agent over the 24-hour window defined by MLE-bench. For conciseness, we will focus our discussion on performance with R1, as the results with o3 follow a similar pattern. We observe that the rankings between agents change over time. For example, at the 3-hour mark, the performance gap between AIRAgreedy and AIRAmcts is notable. This gap narrows by the 10-hour mark, where all agents perform similarly. Meaningful divergence in performance appears only after 19 hours. These trends reflect the interaction between rankings and the resources provided, as shown in Figure 7.
@@ 5.3 The Generalization Gap: Searching with a Proxy Evaluation
Figure 6: a) The validation–test metric mismatch. Results shown are for agents using R1. Bars depict the absolute performance gap between three configurations: (i) using the validation metric for both intermediate (search) and final (submission) node selection; (ii) using the validation metric only for search and the test metric for selection; and (iii) using the test metric for both search and selection. b) Bridging the validation–test gap. Medal rate achieved by two final node selection strategies as a function of $k$: (i) randomly sampling $k$ nodes and reporting the highest test score among them; (ii) selecting the top $k$ nodes by validation score and reporting the highest test score among those. As $k$ increases, the validation‐based strategy closes the gap to the upper‐bound performance given by the best test score over the entire search graph.
@@
How large is the impact of the generalization gap on performance? We begin by comparing two extremes: Val/Val—both intermediate and final node selection use the validation score (standard practice); Test/Test—an oracle baseline using the true test score at both stages. As shown in Fig.˜6, searching based on the test score instead of the validation score improves performance by 9.4 % and 12.4 % for AIRAmcts and AIRAevo. The gap is even higher for the agents using a greedy search policy, reaching 15 % for AIRAgreedy and 16.6 % for AIDEgreedy. For the rest of this section, we investigate the nature of this gap in performance and how to close it.
How much of the gap can be attributed to final node selection? To answer this question, we consider the following two settings: Val/Val; and Val/Test—the intermediate node selection uses the validation score, while the final node selection is made based on the test scores. The results (Fig.˜6) indicate that by selecting the best node in a search graph constructed without privileged access, i.e., based on validation signal, all agents achieve a performance boost of 9 to 13 absolute points. Crucially, comparing Val/Test with Test/Test shows that oracle final-node selection eliminates the gap between the standard (Val/Val) and oracle (Test/Test) settings for both AIRAmcts and AIRAevo. Even for the agents implementing the simpler greedy policy, selecting the best node in the final search graph closes more than 60 % of the gap. These findings highlight robust final-node-selection strategies as a promising path to higher performance.
Bridging the gap through multiple submissions. A straightforward remedy is to hedge against validation score noise by selecting multiple promising nodes. Specifically, among the top-$k$ nodes ranked by validation score in the search graph, we report the test score of the best-performing one. As a baseline, we repeat the process for a set of random-$k$ nodes, and summarize the results in Fig.˜6. We highlight the following observations: 1 the larger gap between top-k and random-k in agents using non-greedy algorithms, AIRAevo and AIRAmcts, suggest greater diversity in their search graphs; 2 with as little as 3 top-k submissions agents achieve an additional 10 % of performance; 3 the top validation nodes are informative of the best performing nodes.